\int \left(x^3+1\right)^5x^2dx
1
Expert's answer
2020-10-05T16:48:24-0400

"\int \left(x^3+1\right)^5x^2dx"


Here, we apply substitution method I.e.


Let "u = x^3 + 1"


"\implies \frac{du}{dx} = 3x^2"


"\implies dx= \frac{1}{3x^2}du"


Replacing back, we have;


"\int (u)^5 \cancel{x^2} \frac{1}{3 \cancel{x^2} }du" "= \frac{1}{3} \int u^5 du"


Now we solve;


"\int u^5 du"


Here, we apply power rule;


Where we let;


"\int u^n du = \frac{u^{n+1}}{n+1}" With "n = 5" , we have;


"= \frac{u^6}{6}"


Replacing this back to "\frac{1}{3} \int u^5 du" we have;


"\frac{u^6}{18}"


Now we undo the substitution

"u = x^3 +1"


"\frac{(x^3 +1)^6}{18}" and we add a C to this, hence the final answer i.e.


"\frac{(x^3 +1)^6}{18} + C"